\begin{thebibliography}{20}

\bibitem{am} H. Abels and A. Manoussos, \emph{Topological generators of abelian Lie groups and hypercyclic finitely generated abelian semigroups of matrices}, Adv. Math. 229 (2012), no. 3, 1862--1872.

\bibitem{AV} H. Abels and E.B. Vinberg, \emph{Generating semisimple groups by tori}, J. Algebra 328 (2011), no 1, 114--121.

\bibitem{Au} H. Auerbach, \emph{Sur les groupes lin'{e}aires born'{e}s (III)}, Studia Math. 5 (1935), 43--49.
\bibitem{BG} E. Breuillard and T. Gelander, \emph{On dense free subgroups of Lie groups}, J. Algebra 261 (2003), no. 2, 448--467.
\bibitem{cao}W. Cao, \emph{Discrete and dense subgroups acting on complex hyperbolic space}, Bull. Aust. Math. Soc. 78 (2008), no. 2, 211--224.


\bibitem{F}N.S. Feldman, \emph{Hypercyclic tuples of operators and somewhere dense orbits}, J. Math. Anal.
Appl. 346 (2008), 82--98.

\bibitem{J2}M. Javaheri, \emph{Semigroups of matrices with dense orbits,} Dyn. Syst. 26 (3) (2011), 235--243.

\bibitem{J}M. Javaheri, \emph{Maximally transitive semigroups of $n \times n$ matrices}, J. Math. Anal. Appl. 401 (2013), 743--753. 


\bibitem{Ku} M. Kuranishi, \emph{On everywhere dense imbedding of free groups in Lie groups}, Nagoya Math. J. 2 (1951), 63--71. 


\bibitem{xia} X. Wang, \emph{Dense subgroups of $n$-dimensional M\"obius groups}, math. Z. 243 (2003), no. 4, 643--651.


\end{thebibliography}

\documentclass[a4paper,12pt]{article}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{amsmath}
\usepackage{latexsym}
\usepackage{epsfig}
\usepackage{amscd}
\usepackage{graphicx}
\usepackage[dvips]{color}
\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}

\theoremstyle{definition}

\theoremstyle{remark}

\title{Dense semigroups of triangular matrices}

\author{ Mohammad Javaheri  \\ Department of Mathematics \\ Siena College, School of Science \\ 515 Loudon Road, Loudonville, NY 12211
\\ \small{Email: mjavaheri@siena.edu}  
}

\begin{document}
\maketitle

\begin{abstract}
Let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$, and $T_n(\mathbb{K})$ be the set of $n\times n$ lower triangular matrices with entries in $\mathbb{K}$. We show that $T_n(\mathbb{K})$ has dense subsemigroups that are generated by $n+1$ matrices.  \end{abstract}

{\small {\emph{2010 Mathematics Subject Classification: Primary 47D03; Secondary 20H20.} 

\emph{Key words and phrases:} Dense subsemigroups, Triangular matrices.}}

\section{Introduction}
Dense subgroups of Lie groups have been studied by many mathematicians, dating back to Auerbach, who proved that every compact semisimple Lie group has a 2-generator dense subgroup \cite{Au}. Kuranishi \cite{Ku} proved that if $G$ is a semisimple Lie group and $a,b$ are two elements near the identity, then the subgroup generated by $a$ and $b$ is dense in $G$ if and only if $\log a$ and $\log b$ generate the Lie algebra $\cal G$. More recently, Abels and Vinberg \cite{AV} showed that every connected semisimple Lie group with finite center has 2-generator dense subsemigroups. In the real case, Breuillard and Gelander \cite{BG} proved that every dense subgroup of a connected semisimple real Lie group has a 2-generator dense subgroup.

Dense subgroups of linear spaces in particular have also been studied. Wang \cite{xia} has shown that every dense subgroup of the group of orientation preserving M\"obius transformations on the $n$-dimensional unit sphere has a dense subgroup generated by at most $n$ elements, where $n\geq 2$. A similar statement was proved by Cao in \cite{cao} for $U(n,1)$. Finally, examples of 2-generator dense subsemigroups of $n\times n$ matrices in both real and complex cases can be constructed \cite{J}.

In connection to hypercyclicity, the following results have inspired the results of this paper. Feldman proved that there exists a dense subsemigroup of $n \times n$ diagonal matrices generated by $n+1$ diagonal matrices \cite{F}. Abels and Manoussos in \cite{am} have shown that the minimum number of generators of a triangular non-diagonalizable abelian subsemigroup with a dense or somewhere dense orbit is $n+1$ in the real case and $n+2$ in the complex case. 

The following theorem is the main result of this article. 

\begin{theorem}\label{maint}
There exist $(n+1)$-generator dense subsemigroups of $n\times n$ triangular matrices in both real and complex cases. 
\end{theorem}

Note that $n+1$ is the least number of generators of a dense subsemigroup of diagonal matrices, and so it is also the least number of generators of a dense subsemigroup of lower triangular matrices.

\section{Proof of the main theorem}
We use Feldman's construction \cite{F} of a hypercyclic $(n+1)$-tuple of $n\times n$ diagonal matrices to construct an $(n+1)$-generator dense subsemigroup of lower triangular matrices. A set of $n\times n$ commuting matrices $(A_1,\ldots, A_k)$ with entries in $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ is called a hypercyclic $k$-tuple, if there exists a vector $X\in \mathbb{K}^n$ such that the set 
$$\left \{A_1^{m_1}\cdots A_k^{m_k}X:m_1,\ldots,m_k \in \mathbb{N} \right \}$$
is dense in $\mathbb{K}^n$. It is straightforward to show that if $(D_0,\ldots, D_n)$ is a hypercyclic $(n+1)$-tuple of diagonal matrices, then the semigroup generated by $D_0,\ldots, D_n$, denoted by $\langle D_0,\ldots, D_n\rangle$ is dense in the set of all $n\times n$ diagonal matrices. In addition, even though Feldman's theorem is stated for $\mathbb{C}$, but the real case can be concluded easily from the complex case. 

In the sequel, all of the statements hold for matrices with entries in $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. Therefore, we might drop the reference to the filed under consideration. We begin with the following lemmas.

\begin{lemma} \label{bounds}
Let $A$ be a lower triangular $n\times n$ matrix with diagonal entries $A_{ii}=a_i$, $1\leq i \leq n$, such that
$$0<|a_1|<\cdots<|a_n|.$$ Then there exists $\lambda>0$, which depends only on $A$, such that for all $1 \leq i,j \leq n$ and all $k\geq 1$, we have
\begin{equation}\label{Alambdaineq}
\left |A^k \right |_{ij} \leq \lambda \left |a_i \right |^k.
\end{equation}
\end{lemma}
The proof of Lemma \ref{bounds} can be found in \cite{J2}. 

Next, we define a total order on the set $\Delta=\{(i,j): 1\leq j \leq i \leq n\}$ as follows:
$$(i,j) \preceq (r,s) \Longleftrightarrow i-j<r-s~\vee~ (i-j=r-s~ \wedge~i \leq r).$$
 Let ${\cal T}_{(r,s)}$ be the set of $n\times n$ lower triangular matrices $T$ such that $T_{ij}=0$ for all $(i,j) \prec (r,s)$. Clearly, if $(r,s) \preceq (t,u)$, then ${\cal T}_{(u,v)} \subseteq {\cal T}_{(r,s)}$. Moreover, each ${\mathcal T}_{(r,s)}$ is closed under matrix multiplication, $(r,s) \in \Delta$.

\begin{lemma}\label{lemars}
Let $D=(a_1,\ldots, a_n)$ be a diagonal matrix such that 
\begin{equation}\label{ineqdiag}
0<|a_1|<\cdots < |a_n|.
\end{equation}
Let $(r,s) \in \Delta$ and $T\in {\cal T}_{(r,s)}$. If $A=D+T$, then
\begin{equation}\label{Akrs}
\left (A^k \right )_{rs}=\left ( \dfrac{a_r^k-a_s^k}{a_r-a_s} \right ) T_{rs}.
\end{equation}
\end{lemma}

\begin{proof}
For $T_1,T_2 \in {\cal T}_{(r,s)}$ and diagonal matrices $D_1$ and $D_2$, one has
\begin{eqnarray}\nonumber
\left [ (D_1+T_1)(D_2+T_2) \right ]_{rs}&=&\sum_{l=1}^n (D_1+T_1)_{rl}(D_2+T_2)_{ls} \\ \nonumber
&=& \sum_{l=s}^r (D_1+T_1)_{rl}(D_2+T_2)_{ls}\\ \nonumber
&=& (D_1+T_1)_{rs}(D_2+T_2)_{ss}+(D_1+T_1)_{rr}(D_2+T_2)_{rs}\\ \label{Akrs2}
&=& (T_1)_{rs}(D_1)_{ss}+(D_2)_{rr}(T_2)_{rs},
\end{eqnarray}
since for $s<l<r$, we have $(r,l) \preceq (r,s)$, and so $(D_1+T_1)_{rl}=(T_1)_{rl}=0$. The equation \eqref{Akrs} then follows from \eqref{Akrs2} and induction on $k$. 
\end{proof}

\begin{lemma}\label{alldiag}
Let $\langle D_\alpha: \alpha \in J \cup \{0\} \rangle$ be a dense subsemigroup of diagonal matrices, where $J$ is a finite index set. Suppose that the diagonal matrix $D_0=(a_1,\ldots, a_n)$ is such that \eqref{ineqdiag} holds. Let $T$ be a lower triangular matrix with $T_{ii}=0$ for all $1\leq i \leq n$. Then the closure of the semigroup generated by $\langle D_0+T, D_\alpha: \alpha \in J \rangle$ contains all diagonal matrices. 
\end{lemma}

\begin{proof}
We prove that if $T\in {\cal T}_{(r,s)}$, where $(r,s) \succeq (2,1)$, then the closure of $\langle D_0+T,D_\alpha: \alpha \in J \rangle$ includes $D_0+T^\prime$ for some $T^\prime \in {\cal T}_{(r,s)^\prime}$, where $(r,s)^\prime$ is the successor of $(r,s)$ in the ordered set $(\Delta,\preceq)$. Equivalently, we show that there exists $T^\prime \in {\cal T}_{(r,s)}$ such that $D_0+T^\prime$ belongs to the closure of $\langle D_0+T,D_\alpha: \alpha \in J \rangle$ and in addition $(T^\prime)_{rs}=0$. It will then follow from a finite induction on the ordered set $(\Delta,\preceq)$, starting with $(2,1) \in \Delta$ and ending in $(n,1) \in \Delta$, that $D_0$ belongs to the closure of $\langle D_0+T,D_\alpha: \alpha \in J \rangle$, and the claim follows from the assumption that $\langle D_\alpha: \alpha \in J \cup \{0\} \rangle$ is dense in the set of diagonal matrices. 

Let $A=D_0+T$. By our density assumption, for each $\alpha \in J\cup \{0\}$, there exists a sequence of exponents $\{d_\alpha^k\}_{k=1}^\infty$ such that $\lim_{k\rightarrow \infty} d_\alpha^k =\infty$ and 
\begin{equation}\label{defB}
\left ( \prod_{\alpha \in J} D_\alpha^{d_\alpha^k} \right )D_0^{d_0^k} \rightarrow B,
\end{equation}
as $k\rightarrow \infty$, where $B$ is an arbitrary diagonal matrix. Let
$$M_k=\left ( \prod_{\alpha \in J} D_\alpha^{d_\alpha^k} \right )A^{d_0^k}.$$
It follows from \eqref{defB} that for all $1 \leq i \leq n$, we have
$$\lim_{k \rightarrow \infty}(M_k)_{ii}=B_{ii}.$$
Since ${\cal T}_{(r,s)}$ is closed under multiplication, we have $M_k-B \in {\cal T}_{(r,s)}$. Moreover, it follows from Lemma \ref{bounds} that for all $1\leq j \leq i$:
\begin{eqnarray} \nonumber
\left | M_k \right |  _{ij}&= & \left |\left ( \prod_{\alpha \in J} D_\alpha^{d_\alpha^k} \right )A^{d_0^k} \right |_{ij} \\ \nonumber
& = &\left |  \prod_{\alpha \in J} D_\alpha^{d_\alpha^k}  \right |_{ii} \cdot \left |A^{d_0^k} \right |_{ij} \\ \nonumber
& \leq &  \lambda \left |  \prod_{\alpha \in J} D_\alpha^{d_\alpha^k}  \right |_{ii} \cdot \left |D_0^{d_0^k} \right |_{ii}
 \leq  \lambda \left | \left ( \prod_{\alpha \in J} D_\alpha^{d_\alpha^k} \right ) D_0^{d_0^k} \right |_{ii} \rightarrow \lambda |B_{ii}|,
\end{eqnarray}
as $k \rightarrow \infty$. In particular, for each $(i,j)$ with $1\leq i,j \leq n$, the sequence $\{(M_k)_{ij}\}_{k=1}^\infty$ is a bounded sequence, hence, by deriving a common convergent subsequence, we obtain a matrix $M$ in the closure of $\langle D_0+T,D_\alpha: \alpha \in J \rangle$ such that $M-B \in {\cal T}_{(r,s)}$. 

Next, by Lemma \ref{lemars}, we have
\begin{eqnarray}\nonumber
(M_k)_{rs} & = & \left [\left ( \prod_{\alpha \in J} D_\alpha^{d_\alpha^k} \right )A^{d_0^k} \right ]_{rs} \\ \nonumber
&=&\left ( \prod_{\alpha \in J} D_\alpha^{d_\alpha^k} \right )_{rr} \left (A^{d_0^k} \right )_{rs} \\ \nonumber
& =& \left ( \prod_{\alpha \in J} D_\alpha^{d_\alpha^k} \right )_{rr} \left (\dfrac{a_r^{d_0^k}-a_s^{d_0^k}}{a_r-a_s} \right ) T_{rs}\\ \nonumber
&=& \left [\left ( \prod_{\alpha \in J} D_\alpha^{d_\alpha^k} \right ) D_0^{d_0^k}\right ]_{rr}\left (\dfrac{1-(a_s/a_r)^{d_0^k}}{a_r-a_s} \right ) T_{rs}\\ \nonumber
& \rightarrow & \dfrac{B_{rr}T_{rs}}{a_r-a_s},
\end{eqnarray}
as $k \rightarrow \infty$, since $r>s$, and so $|a_s/a_r|<1$. 
It follows that
$$M_{rs}= \lim_{l \rightarrow \infty} (M_{k})_{rs}= \dfrac{B_{rr}T_{rs}}{a_r-a_s}.$$
So far, we have proved that for any arbitrary diagonal matrix $B$, there exists $M=\eta(B)$ in the closure of $\langle D_0+T,D_\alpha: \alpha \in J\rangle$ such that $M-B \in {\cal T}_{(r,s)}$ and $M_{rs}=B_{rr}T_{rs}/(a_r-a_s)$. Let $B_2$ be the diagonal matrix with $(B_2)_{rr}=-1$ and $(B_2)_{ii}=1$ for all $i\neq r$. Also, let $B_1=B_2D_0$. Finally, let $M_1=\eta(B_1)$ and $M_2=\eta(B_2)$. Then for $T^\prime=M_1M_2-D_0$, we have $T^\prime \in {\cal T}_{(r,s)}$ and $D_0+T^\prime=M_1M_2$, which belongs to the closure of $\langle D_0+T,D_\alpha: \alpha \in J \rangle$. Moreover, 
\begin{eqnarray}\nonumber
\left ( T^\prime  \right )_{rs}=(M_1M_2)_{rs}&=&\sum_{i=1}^n (M_1)_{ri}(M_2)_{is}\\ \nonumber
&=& (M_1)_{rs}(M_2)_{ss}+(M_1)_{rr}(M_2)_{rs} \\ \nonumber
&=& \dfrac{(B_1)_{rr}T_{rs}}{a_r-a_s} (B_2)_{ss}+(B_1)_{rr} \dfrac{(B_2)_{rr}T_{rs}}{a_r-a_s} \\ \nonumber
&=& \dfrac{-a_rT_{rs}}{a_r-a_s}+\dfrac{a_r T_{rs}}{a_r-a_s}=0,
\end{eqnarray}
and the claim follows. 
 \end{proof}

\begin{theorem}\label{main}
Suppose that the semigroup $\langle D_\alpha: \alpha \in J\cup \{0\} \rangle$ of $n\times n$ diagonal matrices is dense in the set of all $n\times n$ diagonal matrices with entries in $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$, where $J$ is a finite index set. Suppose that $D_0=(a_1,\ldots, a_n)$ is such that the inequalities in \eqref{ineqdiag} hold. Let $T$ be a matrix such that 
\begin{equation}\label{condtr}
T_{ij} \neq 0 \Longleftrightarrow 1 \leq j <i \leq n.
\end{equation}
 Then $\langle D_0+T,D_\alpha: \alpha \in J\rangle$ is dense in the set of all $n\times n$ lower triangular matrices with entries in $\mathbb{K}$. 
\end{theorem}

\begin{proof}By Lemma \ref{alldiag}, the closure of $\langle D_0+T,D_\alpha: \alpha \in J\rangle$ contains all diagonal matrices. 

Proof of Theorem \ref{main} is by induction on $n$. The case $n=1$ is trivial. Therefore, suppose the claim is true for $n-1$, where $n\geq 2$. Given an $n\times n$ matrix $X$, let $\sigma(X)$ be the upper left $(n-1)\times (n-1)$ block of $X$. In particular, the semigroup generated by $\langle \sigma(D_\alpha): \alpha\in J \cup \{0\} \rangle$ is dense in the set of diagonal $(n-1)\times (n-1)$ matrices. It follows from the inductive hypothesis that 
$\langle \sigma(D_0+T), \sigma(D_\alpha): \alpha \in J \rangle$
is dense in the set of $(n-1)\times (n-1)$ lower triangular matrices. In particular, for each lower triangular $(n-1)\times (n-1)$ matrix $R$, there exists a sequence 
$$L_i=\begin{bmatrix} R_i & 0 \\ V_i & x_i\end{bmatrix} \in \bar \Lambda ,~i\geq 1,$$ such that $R_i \rightarrow R$ as $i\rightarrow \infty$, and $x_i \neq 0$. Here $\bar \Lambda$ is the closure of $\Lambda=\langle D_0+T,D_\alpha: \alpha \in J \rangle$. Choose a sequence of nonzero numbers $\{a_i\}_{i=1}^\infty$ such that $\lim_{i\rightarrow \infty}a_iV_i=0$. Then, we have
\begin{equation}\label{mult3}
\begin{bmatrix} I_{n-1} & 0 \\ 0 & a_i \end{bmatrix} L_i \begin{bmatrix} I_{n-1} & 0 \\ 0 & x/(a_ix_i) \end{bmatrix}=\begin{bmatrix} R_i & 0 \\ a_iV_i & x \end{bmatrix} \rightarrow \begin{bmatrix} R & 0 \\ {\bf 0} & x \end{bmatrix}, 
\end{equation}
as $i \rightarrow \infty$, which implies that: 
\begin{equation}\label{limmat}
\begin{bmatrix}R & 0 \\ {\bf 0} & x \end{bmatrix} \in \bar \Lambda,
\end{equation}
for all $(n-1)\times (n-1)$ lower triangular matrices and any given $x\in \mathbb{R}$, since the three matrices on the left side of equation \eqref{mult3} belong to $\bar \Lambda$ (note that all diagonal matrices belong to $\bar \Lambda$ by Lemma \ref{alldiag}). 
Next, we show that every $n\times n$ triangular matrix is a product of two matrices of the form \eqref{limmat} with $D_0+T$. We write
$$D_0+T=\begin{bmatrix}A^\prime & 0 \\ {V} & a_n \end{bmatrix}.$$
Let 
$$B=\begin{bmatrix}U & 0 \\ {W} & z \end{bmatrix}$$ 
be any lower triangular matrix such that vector $W$ has no zero entries, and $z \neq 0$. Let $x={z}/{a_n}$ and choose an invertible diagonal matrix $S$ such that $xVS=W$. Finally, let $R=US^{-1}(A^\prime)^{-1}$. One has
$$\begin{bmatrix} R & 0 \\ {\bf 0} & x \end{bmatrix}  \begin{bmatrix}A^\prime & 0 \\ {V} & a_n \end{bmatrix} \begin{bmatrix} S & 0 \\ {\bf 0} & 1 \end{bmatrix}=\begin{bmatrix} RA^\prime S & 0 \\ xVS & xa_n \end{bmatrix}=\begin{bmatrix}U & 0 \\ {W} & z \end{bmatrix},$$
which implies that $B \in \bar \Lambda$. Since $\bar \Lambda$ is closed, the conditions that $W$ has no zero entries and $z\neq 0$ can be removed to conclude that $\bar \Lambda$ contains all lower triangular matrices, and the proof is completed.
\end{proof}

We are now ready to state the proof of Theorem \ref{maint}. 
\\
\\
\noindent \emph{Proof of Theorem \ref{maint}.}
It follows from \cite[Theorem 3.4]{F} and its proof that there exists a dense subsemigroup of diagonal matrices generated by $n+1$ diagonal matrices $D_0,\ldots, D_n$, where the diagonal entries of $D_0=(a_1,\ldots, a_n)$ are such that $|a_i| \neq |a_j|$ for all $i\neq j$. Via a permutation, we can assume that, in addition, the entries of $D_0$ satisfy the inequalities in \eqref{ineqdiag}. It then follows from Theorem \ref{main} that for any matrix $T$ that satisfies condition \eqref{condtr}, the semigroup generated by $D_0+T,D_1,\ldots, D_n$ is dense in the set of $n\times n$ traingular matrices. 
\hfill $\square$

\end{document}